solve x^2-7x-3=0 for x.
Answer by checkley75(3666)

X^2-7X-3=0 USING THE QUADRATIC EQUATION WE GET:
X=(7+-SQRT[-7^2-4*1*-3])/2*1
X=(7+-SQRT[49+12])/2
X=(7+-SQRT61)/2
X=(7+-7.81)/2
X=(7+7.81)/2
X=14.81/2
X=7.405 ANSWER.
X=(7-7.81)/2
X=-.81/2
X=-.405 ANSWER.